Circular No. 7522

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

LEONID METEORS 2000 M. Gyssens, International Meteor Organization, reports that there may

have been several peaks in terms of the intensity of Leonid meteors on Nov. 17 and 18, as seen in preliminary data contributed by numerous observers worldwide. There may have been minor peaks (ZHR ∼ 100) near Nov. 17.25 and 17.33 UT, before activity began rising noticeably around Nov. 18.0. Two significant peaks, during Nov. 18.1–18.2 (ZHR ∼ 200–300) and ∼ 18.30 ± 0.01 (ZHR ∼ 400), appear to confirm the predictions by McNaught and Asher (1999, WGN 27, 85), which are based on the theory that the observed peaks represent debris ejected from comet 55P/Tempel-Tuttle eight and four revolutions ago, respectively.

HERCULES X-1 B. Boroson, College of Wooster; S. D. Vrtilek, Smithsonian Astrophys-

ical Observatory; P. Boyd, NASA Goddard Space Flight Center; and A. Levine, Center for Space Research, Massachusetts Institute of Technology, write: “Data from the All Sky Monitor (ASM) on the Rossi X-ray Timing Explorer (RXTE) suggest that the x-ray binary pulsar Hercules X-1 has emerged from the x-ray low state it entered in Mar. 1999 (IAUC 7139). In the period from Nov. 13.0 to 17.2 UT, four out of five averages of 25 ASM observations each detect Hercules X-1 with at least 15σ significance. The count rate peaks near 5 counts/s, corresponding to a flux of 60 mCrab. The average of 25 observations that was not 15σ-significant overlapped the expected eclipse of Hercules X-1. Hercules X-1 was also detected by the ASM over an ∼ 5-day period with a peak of 6 counts/s (80 mCrab) on Oct. 12 (35 days prior to the current detection). Hercules X-1 historically has shown a 35-day x-ray cycle peaking near 6 counts/s in the RXTE ASM. Pointed x-ray observations to determine the spectrum and spin evolution are encouraged, as are observations at other wavelengths. Real-time

quick-look data can be obtained at http://heasarc.gsfc.nasa.gov/cgi-
bin/xte weather/xte lc?source=HERX1, and daily averages of quick-look

data over the entire mission can be obtained at http://xte.mit.edu/.”

COMET C/1999 T1 (McNAUGHT-HARTLEY) Total visual magnitude estimates: Oct. 27.75 UT, 8.9 (M. Mattiazzo,

Wallaroo, S. Australia, 25×100 binoculars); Nov. 3.72, 8.6 (D. A. J. Seargent, The Entrance, N.S.W., 25×100 binoculars); 19.63, 8.5 (S. T. Rae, Whakatane, New Zealand, 0.15-m reflector).

2000 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green